$20,000 is invested at 16% per year for 3 months. How much interest is earned on the investment?
I've tried 16% divided by 12 = 1 4/12 *3 = 2 1/3 is that right?

Found 2 solutions by rfer, nyc_function:
Answer by rfer(16322)

A=20000(1+.16/12)^3
A=20000(1+0.01333)^3
A=20000*1.04
A=20800
20800-20000=$800

Answer by nyc_function(2741)

I = PRT is the formula you need.
NOTE: T = time (in years)
3 months = 3/12 or 1/4 years
We are searching for I.
I = (20,000)(0.16)(1/4)
I = $800